Ground beef sells for $4.75 per kg and ground pork sells for $ 5.50 per kg. How many kilograms of ground pork should be mixed with 8 kg of ground beef to make a mixture that sells for $5.10 per kg?
Answer by stanbon(75887)

Ground beef sells for $4.75 per kg and ground pork sells for $ 5.50 per kg. How many kilograms of ground pork should be mixed with 8 kg of ground beef to make a mixture that sells for $5.10 per kg?
------------------------------------
Equation:
Quantity Eq:: p + b = 8 kg
Value Eq:::5.5p + 4.75b = 5.1*8 dollars
--------------------------------------------
Multiply thru the Quantity Eq. by 55
Multiply thru the Value Eq. by 10
--------------------------------------------
55p + 55b = 55*8
55p + 47.5b = 51*8
----------------------------
Subtract and solve for "p":
7.5p = 4*8
p = 4.267 kg (amt. of pork needed in the mixture)
---
Solve for "b":
b + p = 8
b + 4.267 = 8
b = 3.733 kg (amt. of beef needed in the mixture)
===================================================
Cheers,
Stan H.